OpenStudy (anonymous):

how do i get the Y in this problem by itself -3x+4y=3

OpenStudy (the_fizicx99):

add 3x to both sides, then you have 4y = 3x + 3 divide everything by 4 y = 3/4x + .75

OpenStudy (♪chibiterasu):

First we have to move the -3x to the other side, so let's add 3x on both sides: -3x + 3x + 4y = 3 4y = 3 + 3x now divide 4 on both sides: 4y/4 = (3 + 3x)/4 y = (3 + 3x)/4

OpenStudy (♪chibiterasu):

Yeah pretty much what he said LOL

OpenStudy (the_fizicx99):

^ I think that wrong.

OpenStudy (♪chibiterasu):

putting it in parenthesis is the same as what you did, it would divide both by 4 so it's really that same.

OpenStudy (♪chibiterasu):

the same*

OpenStudy (the_fizicx99):

The 3/4x would be in that form incase this was y = mx + b Otherwise yes nvm

OpenStudy (the_fizicx99):

for rise/run

OpenStudy (♪chibiterasu):

^^ that :P

OpenStudy (anonymous):

thank you so much algerbra 1 is so hard for me

OpenStudy (anonymous):

this is the rest

OpenStudy (♪chibiterasu):

We're not going to do your entire test. >_>